OpenStudy (anonymous):

GRAPH THIS EQUATION BY USING 3 POINTS-3X+2Y=2

OpenStudy (anonymous):

3x + 2y = 2 we find points for drawing a line putting x = 0, we get 2y =2 or y = 1 point becomes (0, 1) Putting x = 2 we get 6 + 2y = 2 2y = 2-6 = -4 y = -4/2 = -2 point becomes (2,-2) finding third point putting x = -2, we get -6 + 2y = 2 2y = 2+6 = 8 y = 8/2 = 4 third point (-2, 4) Now plot the points (0, 1), (2, -2), (-2, 4) join and get the line

OpenStudy (anonymous):

isn't it this way of plotting (x,y,z)?

OpenStudy (anonymous):

its a 2 dimensional geometry so points will be (x, y)

OpenStudy (anonymous):

aah ok i thought its 3D already cause its possible >.<

OpenStudy (anonymous):

For 3 D we will get equation in 3 variables to solve.